# VibraForge: A Scalable Prototyping Toolkit For Creating Spatialized Vibrotactile Feedback Systems

### v1

@@ 5.3.1. Vibrotactile System Development
The first step of the development was to determine the layout of the vibration units on the body. Since there is no prior work that studied the mapping between tactile feedback positions on the body and human perception of 3D directions, we conducted a perceptual study with 10 participants to identify a mapping. The layout was initialized with 46 vibration units uniformly distributed on the body and optimized using user data. The final layout used 32 vibration units to achieve comprehensive coverage of the 3D space (Figure 16(b)). More details on the perceptual study and layout optimization can be found in (Huang et al., 2024b).
Figure 16. (a) The vibrotactile device designed during Case Study 3 assists UAV operators with collision avoidance by delivering obstacle directions via multi-point vibrotactile feedback on the body. Operators not only see and feel visible obstacles (yellow), but also perceive obstacles outside their field of view (orange). (b) The layout of vibration units on the upper body. Red dots represent the initial layout with uniform spacing, while blue dots represent the final layout and are mapped to directions in the 3D space.
@@
(a) The vibrotactile device designed during Case Study 3 assists UAV operators with collision avoidance by delivering obstacle directions via multi-point vibrotactile feedback on the body. Operators not only see and feel visible obstacles (yellow), but also perceive obstacles outside their field of view (orange). (b) The layout of vibration units on the upper body. Red dots represent the initial layout with uniform spacing, while blue dots represent the final layout and are mapped to directions in the 3D space.
The second step was to affix control units and vibration units on the garment, which was a compression shirt similar to that used in Case Study 2. Thirty-two VCA vibration units were press-fit to the garment. A large control unit was placed on the operator’s back. The vibration units were grouped into four chains, each covering the front left, front right, back left, and back right.
The third step was to integrate the hardware into the simulated environment, which was built using Microsoft AirSim (Shah et al., 2017) with Unreal Engine 4 and Python APIs. Since the mapping between vibration units and 3D directions were known from the perceptual study, when a potential collision was detected, the system would trigger the unit that had the nearest direction with the colliding object in the space. This avoided the need to reproduce the physical layout of the vibration units in the system. The risks of colliding with surrounding obstacles were computed using control barrier functions (Ames et al., 2019) and scaled to $[0,15]$ to match the actuators’ intensity levels. If the risk of an obstacle was above $0.5$, the vibration unit in the corresponding direction would be activated by sending commands through a Python server.
@@ 5.3.2. User Study Design
For each condition, the experimental scene was a $5m\times 5m\times 50m$ tunnel (Figure 17(a)) that required participants to steer quadrotors forward, right, or upward. Different flying directions were used to vary visual information capacity as obstacles are often outside the camera’s field of view when flying right and upward. The tunnel contained fifteen objects that were randomly placed to avoid learning effects. Based on these conditions, we hypothesize that (a) vibrotactile feedback would enhance teleoperation performance compared to no feedback, and (b) the improvement would be more effective in the right and upward directions than the forward direction.
Figure 17. (a) A first-person view of the simulation environment used in Case Study 3’s study, with randomly positioned obstacles and a top-down view on the right. (b) The results for the total distance travelled, number of collisions, and input disagreement metrics for the feedback conditions (NA, FSC, VSC) and flying directions (forward-FWD, right-R and upward-UP). The error bars represent the standard error of the mean (SEM), $\mathchoice{{}^{\displaystyle*}}{{}^{\textstyle*}}{{}^{\scriptstyle*}}{{}^{\scriptscriptstyle*}}=p<0.05$, $\mathchoice{{}^{\displaystyle**}}{{}^{\textstyle**}}{{}^{\scriptstyle**}}{{}^{\scriptscriptstyle**}}=p<0.01$, and $\mathchoice{{}^{\displaystyle*{*}*}}{{}^{\textstyle*{*}*}}{{}^{\scriptstyle*{*}*}}{{}^{\scriptscriptstyle*{*}*}}=p<0.001$.
[…]
This figure shows an evaluation of a UAV teleoperation system. In (a), a navigation scene demonstrates the obstacle course where participants navigate using different control conditions. In (b), bar graphs display the performance metrics, including total distance, number of collisions, and input disagreement across multiple conditions (NA, FSC, VSC, etc.), highlighting significant differences between the control methods.
@@ 5.3.3. Analysis and Results
We evaluated three metrics: total distance travelled, number of collisions, and input disagreement (i.e., the deviation of user steering commands from the safe input range during flight). Using a two-way repeated measures ANOVA in SPSS, we assessed the effect of feedback condition and flying direction. Post-hoc pairwise comparisons with a Bonferroni correction were conducted for significant results (Figure 17(b)).
The RM-ANOVA determined that feedback condition had a significant effect on total distance travelled ($F(2,22)=3.585,p<0.05$), as did with flying direction ($F(2,22)=17.868,p<0.001$). The interaction between feedback condition and flying direction was also significant ($F(1.957,21.527)=4.916,p<0.01$). The post-hoc analysis of the interaction effect revealed that FSC-R had longer distance than FSC-FWD ($p<0.01$), FSC-UP ($p<0.01$), NA-R ($p<0.01$), and VSC-R ($p<0.05$), while VSC-FWD had a significantly shorter distance than VSC-R ($p<0.01$) and VSC-UP ($p<0.001$).
@@
In summary, VSC caused fewer collisions and less input disagreement compared to NA, and achieved comparable performance with FSC. Additionally, the right and upward directions had more collisions than the forward condition, highlighting the increasing reliance users had on haptic feedback when under limited visual capacity. Thus, our hypotheses were accepted.
@@ 5.3.4. Lessons Learned
The process of using VibraForge to build the teleoperation vibrotactile system demonstrated several benefits. First, the flexible design (DR4) allowed us to iterate on the layout of vibration units rapidly during the perceptual study. From the initial 46 actuator layout to the final 32 actuator layout, we could easily adjust the positions of vibration units on the body and test various spatial mappings. Second, the low latency vibration responses (DR2) also provided instant feedback for participants, which is essential in high speed teleoperation tasks. In addition, the varied intensities (DR2) provided more nuanced information for participants, helping them prioritize their responses towards obstacles with higher risks. Together, these benefits demonstrate the usability of the toolkit for a complex application.
We also encountered several design challenges. One issue was that participants reported that the same intensity level was perceived differently across body areas. Locations with bone contacts were perceived to be stronger, while those with more subcutaneous tissues were perceived to be weaker. This variability may affect feedback precision during critical applications, and require a deeper investigation into how this can be overcome. Another issue was encountered during the layout iteration. Because the final layout was relatively sparse, the original connection wires that were used were too short. Therefore, we had to solder and assemble new wires. This, however, helped us make the toolkit adaptable to more diverse future use cases.
@@ 6. Discussion
In this section, we discussed the implications drawn from case studies, including strengths and limitations of VibraForge. Additionally, from our first-hand experiences in case studies, we synthesized a generalized five-step pipeline for building spatialized vibrotactile systems. Potential applications of VibraForge were highlighted at the end.
@@ 6.1. Strengths of VibraForge
VibraForge aims to provide a ”least resistance path” (Ledo et al., 2018) towards creating spatialized vibrotactile systems, which was demonstrated through the three case studies.
@@
VibraForge offers great scalability. Systems designed in the case studies are of different scales, ranging from sleeve-type phonemic display with 24 vibration units on the forearm, to body-scale system with over 100 vibration units spanning across the upper body. Compared existing toolkits that are often constrained intrinsically to less than 16 actuators, VibraForge’s chain connection method enables designers to easily scale their system to more than 100 vibration units, significantly expanding the potential design space. Furthermore, adding a new unit to an existing system is also as simple as plug-and-play, without the need for any soldering or programming.
VibraForge offers great expressivity. Information of various complexity are adequately conveyed through spatial vibrotactile feedback in the case studies, ranging from simple on/off status for VR collision detection, to continuous risk levels in teleoperation and symbolic phoneme representation in the phonemic display. Compared with prior spatialized vibrotactile systems that are constrained to using ERMs and low-bandwidth control, our toolkit offers various actuator options and provides fine-grained control over vibration characteristics of each actuator. This enables vibration units to reliably reproduce audio waveforms in technical evaluations.
The GUI Editor of VibraForge features usability. In the phonemic display study, it effectively expedited the pattern authoring process by reducing the average pattern authoring time to one minute. Compared with prior design interfaces, our editor highlighted the canvas and the timeline that provided substantial support for multi-actuator pattern design.
@@
VibraForge also features portability and flexibility. Compared with the original phonemic display that was wired to a bulky desktop audio device, components used in our prototype were more portable and easy to wear, such as the small control unit that communicates via Bluetooth. In the UAV teleoperation study, flexibility ensured that vibration units could be frequently adjusted during the layout optimization perceptual study. Compared to using commercial devices that often have fixed feedback positions, VibraForge enables designers to easily adjust the positions of vibrations units, thus enabling rapid design iterations.
@@ 6.2. Limitations of VibraForge
Although VibraForge can enable designers to quickly prototype and evaluate new multi-actuator vibrotactile devices and experiences, the toolkit does not come without limitations. For example, the open-loop control used in the chain connection topology does pose some issues. First, when a control unit sends a vibration command, it is unclear if the command was successfully transmitted to the target vibration unit or not. This is especially important when sending a ”stop” command, because the vibration unit won’t stop vibrating if the commands was missed during the transmission process. Second, when multiple actuators were activated, there was a noticeable voltage drop along the chains. Although the technical evaluation showed that the supply voltage remained in the MCU operating range and actuator voltage range, it would still affect the users’ subjective experience of the vibrations.
@@
By adding extra command receiver pins on the control unit, one could instead use closed-loop control to solve these problems. With such a design, the vibration units and the control unit would form a loop, wherein the last vibration unit of the chain would be connected back to the control unit using a receiver pin. Whenever a vibration unit receives a command for itself, it would not only process the command and trigger the vibration, but also pass a confirmation message to the next unit. When the last unit in the chain receives this confirmation message, it would send the message back to the central unit, thus closing the loop. In cases where commands fail during the transmission process, the central unit would miss the confirmation message and resend the message. A closed loop connection would also reduce the voltage drop along the chain, as both ends of the chain would be connected to the power source. An LTspice simulation shows that an open-loop, eight-unit chain would result in voltage drop from 5 V to 3.1 V, while a closed-loop version would increase the voltage to 4.2 V.
@@
Additionally, while the GUI Editor’s waveform design was advantageous for the phonemic display pattern design, designers would face difficulties when designing more complex spatial patterns. For example, if a designer wanted to create the effect of actuators vibrating from a central point and expanding out in concentric circles, a designer would have to manually design the waveform, assign the waveform to each vibration unit, and specify the start and end time, which is very tedious. A better solution would be to use graphical design metaphors and breakdown spatial patterns to low-level animations of point, line, and shapes. By leveraging phantom sensation illusions (Schneider et al., 2015), designers would only need to draw the shapes and trajectories, and they would be automatically converted to vibration commands for each unit. In this way, complex patterns could be created with ease.
Our case studies also emphasized the need for an automation pipeline to detect the positions of vibration units on the body and convert them to virtual layout that can be later used in various programming environment. We believe this need can be fulfilled with computer vision and motion capture techniques in the future. Using RGB images of the body, computer vision methods could detect the human body outline, segment vibration units from the body, and record the units’ positions. Similarly, retro-reflective markers could be mounted on top of each vibration unit to create a tracking object in the motion capture system. These positions could then be tracked in real-time and loaded into a development environment such as Unity Engine. In this way, designers could focus on the interaction design rather than manually creating the virtual layout.

### v2

@@ 5.3.1. Vibrotactile System Development
Our first step was to determine the layout of the vibration units on the body. Since there is no prior work on mappings between bodily tactile feedback positions and human perceptions of 3D directions, we conducted a perceptual study with 10 participants to identify a mapping. The layout was initialized with 46 vibration units uniformly distributed on the body and optimized using user data. The final layout used 32 vibration units to achieve comprehensive coverage of the 3D space (Figure 17). More details on the perceptual study and layout optimization can be found in (Huang et al., 2024b). Then we affixed press-fit 32 VCA vibration units to the garment, which was a compression shirt similar to Case Study 2. The vibration units were grouped into four chains, covering the front left, front right, back left, and back right. A large control unit was placed on the operator’s back.
Figure 17. The layout of vibration units on the upper body. Red dots represent the initial layout with uniform spacing, while blue dots represent the final layout and are mapped to directions in the 3D space.
@@
The layout of vibration units on the upper body. The initial layout was designed using 46 units with uniform spacing, while the final layout only used 32 units with more sparse spacing.
We then integrated the hardware into a Microsoft AirSim (Shah et al., 2018) simulation environment, built with Unreal Engine 4 and our Python server. Since the mapping between vibration units and 3D directions were known, when a potential collision was detected, the system would trigger the unit that had the nearest direction with the colliding object in the space. The collision risks with obstacles were computed using control barrier functions (Ames et al., 2019) and scaled to $[0,15]$ to match the actuators’ intensity levels. If the risk of an obstacle was above $0.5$, the vibration unit in the corresponding direction would be activated by sending parameterized vibration commands through the Python server.
Figure 18. A first-person view of the simulation environment used in Case Study 3’s study, with randomly positioned obstacles and a top-down view on the right.
[…]
This figure shows a navigation scene demonstrating the obstacle course where participants navigate using different control conditions.
Figure 19. The results for the total distance traveled, number of collisions, and input disagreement metrics for the feedback conditions (NA, FSC, VSC) and flying directions (forward-FWD, right-R, and upward-UP). The error bars represent the standard error of the mean (SEM), $\mathchoice{{}^{\displaystyle*}}{{}^{\textstyle*}}{{}^{\scriptstyle*}}{{}^{\scriptscriptstyle*}}=p<0.05$, $\mathchoice{{}^{\displaystyle**}}{{}^{\textstyle**}}{{}^{\scriptstyle**}}{{}^{\scriptscriptstyle**}}=p<0.01$, and $\mathchoice{{}^{\displaystyle*{*}*}}{{}^{\textstyle*{*}*}}{{}^{\scriptstyle*{*}*}}{{}^{\scriptscriptstyle*{*}*}}=p<0.001$.
[…]
Bar graphs display the performance metrics, including total distance, number of collisions, and input disagreement across multiple conditions (NA, FSC, VSC, etc.), highlighting significant differences between the control methods.
@@ 5.3.2. User Study Design
For each condition, the experimental scene was a $5m\times 5m\times 50m$ tunnel (Figure 18) that required participants to steer quadrotors forward, right, or upward. Different flying directions were used to vary visual information capacity as obstacles are often outside the camera’s field of view when flying right and upward. Fifteen objects were randomly placed in the tunnel to avoid learning effects. Based on these conditions, we hypothesize that (a) vibrotactile feedback would enhance teleoperation performance compared to no feedback, and (b) the improvement would be more effective in the right and upward directions than the forward direction.
@@ 5.3.3. Analysis and Results
We evaluated three metrics: total distance traveled, number of collisions, and input disagreement (i.e., the deviation of user steering commands from the safe input range during flight). Using a two-way repeated measures ANOVA in SPSS, we assessed the effect of feedback condition and flying direction. Post-hoc pairwise comparisons with a Bonferroni correction were conducted for significant results (Figure 19).
The RM-ANOVA determined that feedback condition had a significant effect on total distance traveled ($F(2,22)=3.585,p<0.05$), as did with flying direction ($F(2,22)=17.868,p<0.001$). The interaction between feedback condition and flying direction was also significant ($F(1.957,21.527)=4.916,p<0.01$). The post-hoc analysis of the interaction effect revealed that FSC-R had longer distance than FSC-FWD ($p<0.01$), FSC-UP ($p<0.01$), NA-R ($p<0.01$), and VSC-R ($p<0.05$), while VSC-FWD had a significantly shorter distance than VSC-R ($p<0.01$) and VSC-UP ($p<0.001$).
@@
In summary, VSC caused fewer collisions and less input disagreement than NA, and achieved comparable performance with FSC. Additionally, the right and upward directions had more collisions than the forward condition, highlighting users’ increasing reliance on haptic feedback when visual capacity was limited. Thus, our hypotheses were accepted.
Figure 20. The usability study environment and examples of the diagrams provided to participants during the study.
[…]
The usability study environment, with a laptop, a mouse, a box of hardware components, and a display. Three slides used in the study showing diagrams of the design task. A task overview with a sketch illustration of the tactile sleeve, a sleeve layout guiding the hardware assembly, and two waveform illustrations for the vibration patterns.
@@ 5.3.4. Lessons Learned
Using VibraForge to build the teleoperation vibrotactile system revealed several benefits. The flexible design (DR4) enabled rapid iterations of vibration unit layouts, from an initial 46-actuator setup to a final 32-actuator configuration, allowing for quick testing of various spatial mappings. Low-latency transmission (DR2) provided immediate feedback essential for high-speed tasks, while varied intensities (DR2) offered nuanced information, helping participants prioritize high-risk obstacles. These features highlight the toolkit’s capability for complex applications.
However, design challenges did arise. Participants noted that intensity perception varied across body areas, with stronger sensations near bones and weaker ones in areas with more tissue. Thus, individual calibrations might be needed to ensure feedback consistency. Additionally, the sparse final layout required soldering longer JST wires to replace the original ones. This suggested that VibraForge needs more enhancements to support sparse layout designs.
6. Usability Study
To understand the usability of the toolkit and determine its ”threshold” (i.e., a user’s ability to get started using the toolkit (Ledo et al., 2018; Myers et al., 2000)), we conducted a first-use lab study (Hartmann et al., 2006) with ten participants recruited from a local technology exhibition. Participants included five males and five females with ages ranging from 18 to 40 years old (Mean = 25.6 years, SD = 6.7). Participants had a range of prior experience programming, building physical devices, and building haptic/tactile devices, with some never having done so and others proclaiming themselves as experts. Each study lasted one hour and each participant received \$30 as compensation. The study was approved by our institutional research ethics board.
6.1. Study Design
The study consisted of a design task, a free exploration activity, and a post-study questionnaire and interview. In the design task, participants replicated a simplified version of the phonemic tactile sleeve display from Case Study 1 (Reed et al., 2018): assembling two chains with six vibration units each and designing two phonemes, consonant /k/ and vowel /au/). Participants received a 5-minute tutorial on the toolkit’s hardware components (i.e., control units, vibration units, connector direction, garment attachment, etc.) before assembling the tactile device following a diagram displayed on a computer monitor. After assembly, participants underwent another 5-minute tutorial on the GUI Editor and waveform creation workflow. They then designed and tested the two phonemes using diagrams on the computer monitor.
For the 15-minute free exploration activity, participants could assemble new devices on other garments (e.g., headband, glove, socks, etc), or combine multiple waveforms to produce complex patterns in the GUI Editor. Participants were encouraged to utilize their creativity, so the experimenter did not provide specific guidance but rather served as an assistant helping participants achieve their design goal. The experimenter recorded the creation process and the final outcome from the exploration. After that, participants completed a questionnaire with demographic questions and questions from the System Usability Scale (SUS) (Brooke, 2013). A semi-structured interview then focused on the overall design experience, key areas for improvement, and potential applications of VibraForge.
6.2. Results
All participants successfully completed the design task and free exploration activity. The average completion time for the design task’s hardware assembly was 8 minutes 39 seconds (SD = 1 minute 52 seconds) and 8 minutes 3 seconds (SD = 2 minutes 46 seconds) for the design task’s software waveform design. These completion times were comparable to the expert time reported in Section 5.1.1. Furthermore, an average SUS score of 76.75 was obtained, which is between the 70th and 80th percentile and higher than average SUS scores (Brooke, 2013). These results indicate that the toolkit has a very low ”threshold” for participants.
Overall, participants appreciated the intuitive design of the hardware components and found the assembly process straightforward and intuitive (e.g., P10: ”the hardware was very easy to assemble and it was very intuitive to configure”; P2: ”it feels like LEGO pieces clicking together”). Participants also appreciated that the software was easy to learn and demonstrated good interface design (e.g., P1: ”the software part was nice because it was similar to the designing tools like Canva or Figma”; P6: ”the timeline at the bottom was very helpful for me when I was trying to debug what was going on”).
During the free exploration activity, participants created a variety of haptic devices (Figure 21). For example, P4 designed a glove with four vibration units on the dorsal side of the hand. P7 attached two vibration units on the left and right sides of the heel on a sock to notify no-entry area. P10 designed a glove with three units on the dorsal side and one unit on the palm to ”play” the Queen song ”We Will Rock You” where each note triggered a different unit on the hand. P2 used the GUI Editor to transcribe the Nat King Cole song “On a Bicycle Built for Two” and played it using the frequency and duration of notes on each vibration unit. P8 made attempts to reproducing other phonemes such as /sh/, /ei/ and /p/ from the original paper (Reed et al., 2018) using the GUI editor.
Figure 21. The devices created by participants during the free exploration: a glove with 4 vibration units on the dorsal side (P4), a sock with two units on the left and right side of the heel (P7), and a glove with three units on the dorsal side and one unit on the palm for transcribing songs (P10).
[…]
The devices created by participants during the free exploration activity: a glove with 4 vibration units on the dorsal side (P4), a sock with two units on the left and right side of the heel (P7), and a glove with three units on the dorsal side and one unit on the palm for transcribing songs (P10).
Participants did note some areas for improvement. For the hardware components, participants mentioned that it would be helpful to have more apparent visual feedback on the chain connections (e.g., P2: “I feel like it [chain connector index] would be much more visible if it’s printed on top of it rather than in the middle.; P7: “I didn’t know whether they [vibration units] were connected … I think making [the LED] more visible would be more helpful”. They also suggested other methods to affix the vibration units to garments to make the toolkit more adaptable to other fabrics (e.g., P7: ”This [bottom ring attachment] and a magnet would have been a bit more effective, because you would only need to press it … and it’s gonna have a really strong grip on it for you to move it around”; P4: ”maybe there’s like a lock method so, when you snap on it locks into the sensors [vibration units]”. Based on these feedback, there is thus a need to add visual indicators for hardware connection statuses and explore additional mechanical affixation structures.
Others expressed the need for more versatile functionality in the GUI Editor. For example, P1 wanted to directly manipulate the waveforms (e.g., ”if I want to change the amplitude, for example, I should just drag the top part and then drag down to lower, like direct manipulation … if I want to change the frequency, I should just click and drag the amount to left or right”. Similarly, because participants needed to drag and drop the same waveform onto several vibration units, P6 suggested adding ”multi-select” functions to edit multiple vibration units concurrently. Because the Waveform Visualizer was used to both edit and assign waveforms, some participants were confused about its current status. P5 commented that ”the fact that the designing of a custom wave and assigning it to the node share the same interface is a bit confusing”. These comments suggest that the GUI Editor requires some refinement to ensure that supported actions are more natural and clearly conveyed to users.
Participants also highlighted several potential applications of the toolkit. P1 and P4 noted how VibraForge could be used to aid visually impaired individuals with navigation. P2 and P10 demonstrated how the toolkit could be used to translate music into tactile patterns, whereas P6 suggested using it to enhance VR and gaming immersion by simulating action feedback. P3 and P9 proposed the creation of devices for personalized massages and stress relief. P7 was enthusiastic about using it for safety applications, such as non-visual notifications while running or cycling, particularly in noisy environments. These examples demonstrate VibraForge’s potential usage in diverse domains.
@@ 7. Discussion
Herein we discuss the strengths and limitations of VibraForge summarized from the studies, and suggested potential solutions to address the limitations. Additionally, based on our first-hand developing experiences, we synthesized a generalized five-step pipeline for building spatialized vibrotactile systems.
@@ 7.1. Strengths of VibraForge
VibraForge aims to provide a ”least resistance path” (Ledo et al., 2018) towards creating spatialized vibrotactile systems via bridging the gap between scalability and expressivity. The systems designed in case studies ranged from a sleeve-type phonemic display on the forearm, to a body-scale system spanning the upper body. Compared to existing toolkits that are often constrained to fewer than 16 actuators (Wittchen et al., 2022; Pezent et al., 2020; Dementyev et al., 2021), VibraForge’s chain connection method enables designers to easily scale their system to more than 100 vibration units, significantly expanding potential design spaces. Furthermore, new units can be plug-and-play into existing systems without additional programming, fostering rapid prototyping process. VibraForge also features great expressivity. It rendered tactile information at different complexities, ranging from simple on/off status for VR collision detection, to parameterized risk levels in teleoperation and symbolic phoneme representation in the phonemic display. Compared with previous spatialized vibrotactile systems that were restricted to using ERMs and PWM control (Wang et al., 2021; Sánchez and Bohne, 2015; Lemmens et al., 2009), our toolkit offers various actuator options and provides fine-grained control over vibration characteristics, thus reliably reproducing complex waveforms.
Furthermore, VibraForge demonstrates high usability. Case Study 1 and the usability study indicated that both professional and novice designers benefited from the streamlined hardware assembly and intuitive GUI design. Compared with prior design interfaces (Pezent et al., 2020; Minamizawa et al., 2012), our GUI Editor provided substantial support for multi-actuator pattern design. Professionals using the GUI Editor effectively reduced the average pattern authoring time. Novice users compared the intuitive hardware assembly to the joyful process of ”building LEGOs”. Within an hour, they learned the toolkit basics and utilized it to finish professional design tasks and for creative exploration. This indicated that the toolkit lowered the technical barrier to building spatialized vibrotactile systems and enhanced the overall efficiency.
@@
VibraForge also features portability and flexibility. Compared with the original phonemic display that was wired to a bulky desktop audio device, our components were more portable and easy to wear, such as the small control unit that communicated via Bluetooth. In the UAV teleoperation, flexibility ensured that vibration units could be frequently adjusted during the layout optimization perceptual study. Compared to using commercial devices (bHaptics, 2024) that often have fixed feedback positions, VibraForge enables designers to easily adjust the positions of vibrations units, thus enabling rapid design iterations.
In summary, rather than contributing one incremental technical advancement, our work fuses multiple components—modularity, usability, and portability—into a cohesive vibrotactile prototyping toolkit. This holistic system-level integration enables a larger design space with a higher “ceiling” (shown in the case studies) and lower “threshold” (shown in the usability study), opening new possibilities for both professional designers and novice users to build applications with haptic feedback.
VibraForge could be useful within several domains. For example, In VR social platforms, spatial vibrotactile feedback could help users express their emotions through social touch interactions (Sykownik and Masuch, 2020). Developers using the toolkit can easily adjust the layout of actuators to build personalized devices, in order to avoid discomfort or sensory issues in certain body areas. As another example, VibraForge can be used to build prototypes for tactile perceptual studies, such as those measuring spatial acuity, information transfer, or other perceptual characteristics. For example, Park et al. performed perceptual studies to measure information transfer across five body positions (Park et al., 2023), but the choice of positions were limited by the layout of bHaptics modules (bHaptics, 2024). By using VibraForge, researchers could freely adjust the layout of vibration units on the body and adapt it to any study requirements.
@@ 7.2. Limitations of VibraForge
The toolkit does not come without limitations. The open-loop control used in the chain connection poses some issues. For example, when a control unit sends a vibration command, it can be unclear if the command was successfully transmitted to the target vibration unit. When transmitting a ”stop” command, this is especially important as a vibration unit will not stop vibrating if the commands was missed. Additionally, when multiple actuators are activated there is a noticeable voltage drop along the chains, which would affect users’ experience of the vibrations. By adding extra command receiver pins to the control unit, the vibration and control unit would form a loop, wherein the last vibration unit in a chain would be connected to the control unit using a receiver pin and could propagate a confirmation message back to the central unit. In cases where commands fail during transmission, the central unit could resend the message. A closed loop connection would also reduce the voltage drop along the chain, as both ends of the chain would be connected to the power source. An LTspice simulation showed that an open-loop, eight-unit chain would result in voltage drop from 5 V to 3.1 V, while a closed-loop would increase the voltage to 4.2 V.
@@
While the GUI Editor’s waveform design was advantageous for the phonemic display’s pattern design, if a designer wanted to create an effect where actuators vibrated out from a central point in concentric circles, a designer would have to manually assign the waveform and specify the start and end time for each vibration unit. Instead, graphical design metaphors could be used to break down spatial patterns into low-level animations that would be automatically converted to vibration commands for each unit.
VibraForge was built focusing on dense matrices of vibration units, as this was consistent with prior spatial tactile feedback systems that employed dense actuator arrays (e.g., facial masks (Wang et al., 2021) and bodysuits (bHaptics, 2024; Lemmens et al., 2009; West et al., 2019)). However, Case Study 3 revealed that it was challenging to support the sparse layout of vibration units on the body, as longer wires were needed to connect more distant units, which may impact the overall system reliability. To overcome this limitation, we plan to support the use of multiple control units within the same system so that designers can place control units at various positions on the body to support such sparse layouts.
Our studies also emphasized the need for an automated pipeline to visualize vibration unit positions. Currently, the GUI Editor canvas provides an approximate virtual layout of vibration units and their relative order within each chain. Although participants in our usability study were able to map physical units to their virtual counterparts, a true-to-scale, integrated visualization would reduce cognitive load and foster more rapid design iterations. One potential solution could be to revise PCB designs and add programmable LEDs to vibration units. When units are activated sequentially on each chain, computer vision algorithms could detect the positions of units on the body and their relative orders. Alternatively, ultra-wideband (UWB) tags could be mounted on top of the vibration units. With multiple UWB anchors setup in the environment, each UWB tag could utilize trilateration to find its own position. We are actively exploring these solutions and will incorporate them into the next iteration.
Lastly, VibraForge currently only supports vibrotactile actuators. Recent multimodal haptic toolkits, such as SleeveIO (Shtarbanov et al., 2023) and TactorBots (Zhou et al., 2023), demonstrated the potential of integrating diverse actuator types (e.g., pneumatic, hydraulic, servo-driven) into haptic systems. However, these toolkits often rely on direct connections between modules and controllers, which limit their scalability. In contrast, VibraForge’s chain-connection architecture naturally supports extensions. Adding new actuator modalities would simply require custom PCB drivers and revised data protocols. Data bytes for intensity and frequency could be repurposed to handle parameters such as servo speed and position, Peltier temperatures, mechanical displacements, and so on. Thus, VibraForge could achieve the expressivity possible with other multimodal toolkits while maintaining its scalability.